OpenStudy (anonymous):

find dy/dx using the chain rule for y=u^4, u=x^3+7.

OpenStudy (anonymous):

\[ \frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx} \]Can you find \(y'(u)\) and \(u'(x)\)?

OpenStudy (ivancsc1996):

\[\frac{ dy }{ dx }=\frac{ dy }{ du }\frac{ du }{ dx }=4u ^{3}3x ^{2}=4(x ^{3}+7)^{3}3x ^{2}\]And then you simplify that.

OpenStudy (anonymous):

ohhok! thank you so much!:)